Ann made
$374
for
17
hours of work.
At the same rate, how much would she make for
12
hours of work?

Answer by jim_thompson5910(35256)

374/17 = x/12

374*12 = 17*x

4488 = 17x

17x = 4488

x = 4488/17

x = 264

-------------------------------------------------------

Question: how much would she make for 12 hours of work?

Answer: $264